OpenStudy (anonymous):

PLEASE HELP!!! ill give medal What is the equation of the line that passes through the points (−2, 3) and (2, 7)? x − y = −1 x − y = −2 x − y = − 5 x − y = − 6

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

I messed up

OpenStudy (mathstudent55):

For points \((x_1, y_1)\) and \((x_2, y_2) \), the equation of the line through them is: \(y - y_1 = \dfrac{y_2 - y_1}{x_2 - x_1}( x - x_1) \)

OpenStudy (anonymous):

slope(m) = (y2 - y1) /(x2 - x1) m = (7 - 3) / (2 - (-2) m = 4/(2 + 2) m = 4/4 m = 1 point slope formula : y - y1 = m(x - x1) y - 7 = 1(x - 2) y - 7 = x - 2 y = x - 2 + 7 y = x + 5 <-- your equation

OpenStudy (anonymous):

x − y = −1 x − y = −2 x − y = − 5 x − y = − 6 these are the answer chocies

OpenStudy (anonymous):

@kelliegirl33 would it be the last one?

OpenStudy (anonymous):

y = x + 5 x - y = -5

OpenStudy (anonymous):

0ok one more Select the equations that are parallel and perpendicular to y = one over fourx + 7 and that pass through the point (−8, 9). parallel: y = one over fourx − 11 perpendicular: y = −4x + 23 parallel: y = −4x + 11 perpendicular: y = one over fourx − 4 parallel: y = 4x − 7 perpendicular: y = negative one over fourx − 7 parallel: y = one over fourx + 11 perpendicular: y = −4x − 23

OpenStudy (anonymous):

I am sorry but I got to go now......I do not have time to do this...sorry

OpenStudy (anonymous):

please! I have to! this is my final exam and this is the last question and then I can submit it. PLEAEE

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (mathstudent55):

I'm here. I'll help. In the future, start a new post for each question.

OpenStudy (anonymous):

ok I just did! please help